# Uniqueness of Maximal Inscribed Parabolas and Minimal Circumscribing Horocycles

### v1

@@ Abstract
We prove existence of three unique “exparabolas” to a triangle. Each of these parabolas is internally tangent to one edge and the two other sides. Among all like parabolas, it is characterized by having maximal parameter. We use this result to prove a more general uniqueness statement on maximal parabolas into a convex point set. In similar spirit, we demonstrate uniqueness of minimal enclosing horocycles in hyperbolic geometry, provided the enclosed set is sufficiently small.
@@ 1 Introduction
A typical technique for proving uniqueness is indirect. Assuming existence of two different extremal (minimal or maximal) ellipsoids of equal size, existence of a further circumscribing/inscribing ellipsoid that is smaller/larger is shown [3, 16, 17]. This is clearly exemplified when proving uniqueness of minimal circumscribing spheres (Figure 1). The common interior of two congruent but distinct spheres can be covered by a smaller sphere. The proof is valid for any reasonable function to measure the size of the sphere, since it can be related to a strictly monotone increasing function of the sphere’s radius. It is also valid for spherical geometry and hyperbolic geometry (existence of a circumscribing sphere needs to be assumed though). Note that the maximal inscribed sphere is not at all unique as the convex hull of two different congruent spheres contains infinitely many congruent spheres.
Figure 1: Uniqueness of minimal circumscribing spheres.
Uniqueness theorems have also been shown for minimal hyperbolas that enclose a set of lines in the plane [18] and for minimal enclosing ellipses in the elliptic and in the hyperbolic plane [19, 20]. The proofs are even more difficult and guarantee uniqueness only under some additional assumptions on the set to be enclosed and the set of conics to minimize over.
@@
Classes of congruent parabolas or horocycles are determined by only one invariant (the parameter in case of parabolas), much like the spheres whose size is unquely determined by the radius. This property is very beneficial in proofs of uniqueness and makes the study of different size functions unnecessary. It should be noted that parabolas and horocycles are unbounded but with only one point at infinity. Hence, unlike ellipsoids, they can be used for conservative (worst case) approximations of certain unbounded point sets.
We continue this text be setting notation and introducing some basic concepts in Section 2 before turning to maximal parabolas in Section 3. Our first result (Theorem 1) shows that to any triangle three locally maximal parabolas can be associated. These exparabolas seem to be an interesting new concept of triangle geometry. The uniqueness result of maximal parabolas in more general settings is then dealt with in Theorem 2. Section 4 is dedicated to the study of minimal enclosing horocycles. The main result, Theorem 3, claims their uniqueness among all enclosing horocycles whose axis ratio (when viewed as a euclidean ellipse) does not drop beneath $\sqrt{2}$.
@@ 2.1 The Common Interior of Two Conics
$p^{t}\cdot C_{0}\cdot p<0\quad\text{and}\quad p^{t}\cdot C_{1}\cdot p<0$
implies
$p^{t}\cdot C(t)\cdot p=(1-t)p^{t}\cdot C_{0}\cdot p+tp^{t}\cdot C_{1}\cdot p<0$
In other words, interior points of both $C_{0}$ and $C_{1}$ are also interior points of $C(t)$.
@@ 2.2 Parabolas and their Size
While in affine geometry, any two parabolas are equivalent, this is not the case in Euclidean geometry. In fact, any two parabolas correspond in a Euclidean transformation plus a uniform scaling. Thus, the “size” of a parabola can be measured as scaling factor relative to a designated “unit parabola.” Another way of embodying the same idea uses the parabola’s parameter $p$ for the purpose of measuring and comparing size. The parameter is the distance between the parabola’s directrix and its focal point [22, Chapter 2]. Therefore, a parabola $P_{1}$ with parameter $p_{1}$ can be mapped into the interior of a parabola $P_{2}$ with parameter $p_{2}$ by a Euclidean displacement, if and only if $p_{1}<p_{2}$. As to the respective size of $P_{1}$ and $P_{2}$ this means that
@@ 2.3 Horocycles and their Size
Horocycles are defined as regular projective conics with all but one point in the interior of $\mathcal{N}$. In other words, they are Euclidean ellipses that hyperosculate $\mathcal{N}$ in a minor vertex (Figure 2). Both, parabolas in Euclidean geometry and horocycles in hyperbolic geometry share the property of having precisely one point at infinity, so that, arguably, horocycles can be regarded as the “parabolas of hyperbolic geometry”. This point of view ties together the different results we prove in this paper. It should be noted, however, that other concepts of a parabola in hyperbolic geometry exists, cf. [24, 25].
Figure 2: Some horocycles in the Cayley-Klein model of hyperbolic geometry.
@@
The hyperbolic area of a horocycle is infinite. But, just as in the case of parabolas in Euclidean geometry, we say that horocycle $h_{1}$ is smaller than horocycle $h_{2}$ if $h_{1}$ one fits into $h_{2}$.” Since any two horocycles correspond in a transformation of hyperbolic geometry (a projective transformation that fixed $\mathcal{N}$ and its interior), the need for this comparison of two horocycles suggest to restrict the transformation group to Euclidean symmetries of $\mathcal{N}$. We already did something similar for parabolas, cf. Remark 1. With this notion of “size comparison” we can to measure the size of a horocycle by an arbitrary strictly monotone function $\operatorname{size}(a)$ of its (Euclidean) major axis length $a$. We will simply use $\operatorname{size}(a)=a$ and, by abuse of notation, also write $\operatorname{size}H=a$.
@@
Section 2.2 actually suggests to use the horocycle’s parameter (when viewed as ellipse of the Euclidean plane) to measure its size. This is consistent with our approach: The parameter of a conic equals the curvature radius on the conic’s principal axis [22, Theorem 2.1.5]. For horocycles this value is $a^{3}$ which is, indeed a strictly monotone function of $a$.
@@ 3.1 Exparabolas of a Triangle
Figure 3: Triangle with in- and excircles.
It is well-known that there exist four circles tangent to all three sides of $\triangle ABC$. One of them, the incircle, is contained in the intersection of the triangle’s positive half planes, while three, the excircles, are contained in the intersection of one negative and two positive half-planes (Figure 3). Clearly, incircle and excircles are of maximal size among all circles contained in the respective intersection of three triangle half-planes and tangent to the three triangle sides. This motivates the next definition.
@@
A parabola $P$ is called exparabola of a triangle $\triangle ABC$ if it is
  • 1.
    tangent to all three triangle sides,
  • 2.
    contained in the intersection of one negative and two positive half-planes of the triangle and
  • 3.
    of maximal size (its parameter is maximal) among all these parabolas.
Figure 4: Triangle with three exparabolas $P_{A}$, $P_{B}$, and $P_{C}$.
Figure 4 displays a triangle $\triangle ABC$ together with its three exparabolas. Existence of precisely three exparabolas is the content of the next theorem.
@@
There exist precisely three exparabolas to a triangle $\triangle ABC$, one in each of the three possible intersections of one negative and two positive half-planes.
@@
we see that at least one root lies between $a_{1}$ and $b_{1}$, that is, at least one of these parabolas contained in the intersection of the negative half-space through $A$ and $B$ and the positive half-spaces through the remaining triangle sides. For reasons of symmetry, exactly one parabola has this property and the other two parabolas are contained, respectively, in the other negative half spaces. ∎
@@ 4 Uniqueness of Minimal Horocycles
In this section we prove a uniqueness result for enclosing horocycles of minimal size in hyperbolic geometry. Any compact point set $\mathbb{F}$ in the hyperbolic plane can be enclosed by suitably chosen horocycle and there also exists an enclosing horocycle of minimal size. It turns out that this minimal horocycle is not generally unique but it is under a natural additional assumption on the point set $\mathbb{F}$.
@@
The assumptions of the theorem guarantee existence of an enclosing horocycle of $\mathbb{F}$ with $\operatorname{size}(\mathbb{F})<2^{-1/2}$. Assuming $H_{0}$, $H_{1}$ are both enclosing horocycles of minimal size. Then $\operatorname{size}(H_{0})=\operatorname{size}(H_{1})<2^{-1/2}$. But this is a contradiction to Lemma 1. ∎
@@ 5 Conclusions
We have proved uniqueness results for maximal enclosed parabolas in the Euclidean plane and of minimal circumscribing horocycles in the hyperbolic plane. These results pose the natural question for uniqueness results for minimal circumscribing parabolas and maximal enclosed horocycles. Unfortunately, neither our proof of Theorem 2 nor of Theorem 3 can be easily adapted to the other case. The adaption of the proof of Theorem 2 fails to the non-existence of linear dual pencils or horocycles, the adaption of Theorem 3 fails to the non-uniqueness of a parabola through just one extremal point $L$.
The result (and also the proof) of Theorem 1 on existence of maximal parabolas inscribed into a triangle is needed in order to prove Theorem 2, the general uniqueness result on maximal parabolas. It is, however, of interest in its own right as it introduces the concept of exparabolas to triangle geometry. Their position within modern triangle geometry is yet to be clarified. This is topic of future research.

### v2

@@ Abstract
We prove existence of three unique “max-exparabolas” to a triangle. Each of these parabolas is internally tangent to one edge and the two other sides. Among all like parabolas, it is characterized by having maximal parameter. We use this result to prove a more general uniqueness statement on maximal parabolas in a convex point set. In similar spirit, we demonstrate uniqueness of minimal enclosing horocycles in hyperbolic geometry, provided the enclosed set is sufficiently small.
@@ 1 Introduction
A typical technique for proving uniqueness is indirect. Assuming existence of two different extremal (minimal or maximal) ellipsoids of equal size, existence of a further circumscribing/inscribing ellipsoid that is smaller/larger is shown [3, 16, 17]. Uniqueness theorems have also been proved for minimal hyperbolas that enclose a set of lines in the plane [18] and for minimal enclosing ellipses in the elliptic and in the hyperbolic plane [19, 20]. These proofs are even more difficult and guarantee uniqueness only under some additional assumptions on the set to be enclosed and the set of conics to minimize over.
@@
Classes of congruent parabolas or horocycles are determined by only one invariant (the parameter in case of parabolas), much like the spheres whose size is uniquely determined by the radius. This property is very beneficial in proofs of uniqueness and makes the study of different size functions unnecessary. It should be noted that parabolas and horocycles are unbounded but with only one point at infinity. Hence, unlike ellipsoids, they can be used for conservative (worst case) approximations of certain unbounded point sets.
We continue this text by setting notation and introducing some basic concepts in Section 2 before turning to maximal parabolas in Section 3. Our first result (Theorem 1) shows that to any triangle three locally maximal parabolas can be associated. These max-exparabolas seem to be an interesting new concept of triangle geometry. The uniqueness result of maximal parabolas in more general settings is then dealt with in Theorem 2. Section 4 is dedicated to the study of minimal enclosing horocycles. The main result, Theorem 3, claims their uniqueness among all enclosing horocycles whose axis ratio (when viewed as a Euclidean ellipse) does not drop beneath $\sqrt{2}$.
@@ 2.1 The Common Interior of Two Conics
$p^{t}\cdot C(t)\cdot p=(1-t)p^{t}\cdot C_{0}\cdot p+tp^{t}\cdot C_{1}\cdot p$
so that interior points of both $C_{0}$ and $C_{1}$ are also interior points of $C(t)$.
@@ 2.2 Parabolas and their Size
While in affine geometry, any two parabolas are equivalent, this is not the case in Euclidean geometry. In fact, any two parabolas correspond in a Euclidean transformation plus a uniform scaling. Thus, the “size” of a parabola can be measured as scaling factor relative to a designated “unit parabola.” Another way of embodying the same idea uses the parabola’s parameter $p$ (the distance between focal point and directrix [22, Chapter 2]) for the purpose of measuring and comparing size. A parabola $P_{1}$ with parameter $p_{1}$ can be mapped into the interior of a parabola $P_{2}$ with parameter $p_{2}$ by a Euclidean displacement, if and only if $p_{1}<p_{2}$. As to the respective size of $P_{1}$ and $P_{2}$ this means that
@@ 2.3 Horocycles and their Size
Horocycles are defined as regular projective conics with all but one point in the interior of $\mathcal{N}$. In other words, they are Euclidean ellipses that hyperosculate $\mathcal{N}$ in a minor vertex, cf. Figure 2 for some examples. Both, parabolas in Euclidean geometry and horocycles in hyperbolic geometry share the property of having precisely one point at infinity, so that, arguably, horocycles can be regarded as the “parabolas of hyperbolic geometry”. This point of view ties together the different results we prove in this paper. It should be noted, however, that other concepts of a parabola in hyperbolic geometry exists, cf. [24, 25].
@@
The hyperbolic area of a horocycle is infinite. But, just as in the case of parabolas in Euclidean geometry, we say that horocycle $h_{1}$ is smaller than horocycle $h_{2}$ if $h_{1}$ one fits into $h_{2}$.” Since any two horocycles correspond in a transformation of hyperbolic geometry (a projective transformation that fixed $\mathcal{N}$ and its interior), the need for this comparison of two horocycles suggests to restrict the transformation group to Euclidean symmetries of $\mathcal{N}$. We already did something similar for parabolas, cf. Remark 1. With this notion of “size comparison” we can reasonably measure the size of a horocycle $H$ by the function $\operatorname{size}H=a$ with $a$ being the (Euclidean) major axis length $a$.
@@
Section 2.2 actually suggests to use the horocycle’s parameter (when viewed as ellipse of the Euclidean plane) to measure its size. This is consistent with our approach: The parameter of a conic equals the curvature radius on the conic’s principal axis [22, Theorem 2.1.5]. For horocycles this value is $a^{3}$ which is a strictly monotone function of $a$.
@@ 3.1 Max-Exparabolas of a Triangle
It is well-known that there exist four circles tangent to all three sides of $\triangle ABC$. One of them, the incircle, is contained in the intersection of the triangle’s positive half planes, while three, the excircles, are contained in the intersection of one negative and two positive half-planes. Clearly, incircle and excircles are of maximal size among all circles contained in the respective intersection of three triangle half-planes and tangent to the three triangle sides. This motivates the next definition.
@@
A parabola $P$ is called an exparabola of a triangle $\triangle ABC$ if it is tangent to all three triangle sides. An exparabola is called a max-exparabola of $\triangle ABC$ if it is of maximal size (parameter) among all exparabolas contained in the intersection of one negative and two positive half-planes of the triangle.
Figure 1: Triangle with three max-exparabolas $P_{A}$, $P_{B}$, and $P_{C}$.
Figure 1 displays a triangle $\triangle ABC$ together with its three max-exparabolas. Existence of precisely three max-exparabolas is the content of the next theorem.
@@
There exist precisely three max-exparabolas to a triangle $\triangle ABC$, one in each of the three possible intersections of one negative and two positive half-planes.
@@
we see that at least one root lies between $a_{1}$ and $b_{1}$, that is, at least one of these parabolas is contained in the intersection of the negative half-space through $A$ and $B$ and the positive half-spaces through the remaining triangle sides. For reasons of symmetry, exactly one parabola has this property and the other two parabolas are contained, respectively, in the other negative half spaces. ∎
@@ 4 Uniqueness of Minimal Horocycles
In this section we prove a uniqueness result for enclosing horocycles of minimal size in hyperbolic geometry. Any compact point set $\mathbb{F}$ in the hyperbolic plane can be enclosed by a suitably chosen horocycle and there also exists an enclosing horocycle of minimal size. It turns out that this minimal horocycle is not generally unique but it is under a natural additional assumption on the point set $\mathbb{F}$.
@@
The assumptions of the theorem guarantee existence of an enclosing horocycle of $\mathbb{F}$ with $\operatorname{size}(\mathbb{F})<2^{-1/2}$. Assume that $H_{0}$, $H_{1}$ are both enclosing horocycles of minimal size. Then $\operatorname{size}(H_{0})=\operatorname{size}(H_{1})<2^{-1/2}$. But this is a contradiction to Lemma 1. ∎
@@ 5 Conclusions
We have proved uniqueness results for maximal enclosed parabolas in the Euclidean plane and of minimal circumscribing horocycles in the hyperbolic plane. These results pose the natural question for uniqueness results for minimal circumscribing parabolas and maximal enclosed horocycles. Unfortunately, neither our proof of Theorem 2 nor of Theorem 3 can be easily adapted to the other case. The adaptation of the proof of Theorem 2 fails to the non-existence of linear dual pencils or horocycles, the adaptation of Theorem 3 fails due to the non-uniqueness of a parabola through just one extremal point $L$.
The result (and also the proof) of Theorem 1 on existence of maximal parabolas inscribed into a triangle is needed in order to prove Theorem 2, the general uniqueness result on maximal parabolas. It is, however, of interest in its own right as it introduces the concept of max-exparabolas to triangle geometry. Their position within modern triangle geometry is yet to be clarified. This is topic of future research.